# Feature Distillation is the Better Choice for Model-Heterogeneous Federated Learning

### v1

@@ 1 Introduction
To address this issue, knowledge distillation DBLP:journals/corr/abs-1811-11479 has emerged as a promising approach, focusing on aggregating the output soft predictions of multiple local models into the global model. The authors in wu2022communication propose aggregating the logit output instead of model parameters and li2019fedmd utilize the aggregated class score to regulate the local training process. Building on the partial parameter aggregation proposed in diao2020heterofl , many works use the ensemble distillation to improve the global model like lin2020ensemble ; wang2023dafkd . wang2022knowledge and zhang2023target have focused on knowledge selection with logit distillation to optimize the aggregated model. shao2024selective identifies the accurate and precise knowledge from local and ensemble predictions, enhancing performance and privacy.
@@ 3.3 FedFD: Feature Distillation for Hetero-FL
Instead of maintaining the projection layer separately for each client, the server now only needs to train $(m-1)$ projection layers $\{\mathcal{M}_{2},\ldots,\mathcal{M}_{m}\}$. This not only reduces training parameters but also ensures that each projection layer has sufficient knowledge for distillation.
@@ 4.1 Experiment Setup
Configurations: Unless otherwise mentioned, we set the number of local training epoch E = 10, communication round T = 200, and the client number K = 20 with an active ratio r = 0.4. For local training, the batch size is 64 and the weight decay is $41e-$ 4. The learning rate is $0.01$ for distillation and $0.001$ for training the local model. For the model on the server, we employ ResNet-18 he2016deep as the basic backbone. Like diao2020heterofl , we construct ten different computation complexity levels {a, b, c,…,j} with the hidden channel decay rate 10%. For example, model "a" has all the model parameters, while models "b" and "c" have the 90% and 80% effective parameters. We use "a-d-g" three different model architectures in our experiments and conduct more experiments with other architectures in Section 4.2. Similarly to diao2020heterofl , we have conducted statistics based on the performance of the global model for all secondary experiments.
@@
Classic FL HeteroFL Local 80.11$\pm$0.95 75.83$\pm$1.35 63.53$\pm$4.66 53.37$\pm$1.11 49.33$\pm$1.76 38.07$\pm$4.52 29.71$\pm$2.85 24.12$\pm$5.03 18.54$\pm$3.40
Global 88.45$\pm$0.03 87.53$\pm$0.15 78.02$\pm$0.65 58.51$\pm$0.19 57.42$\pm$0.12 53.98$\pm$0.43 32.38$\pm$0.51 29.88$\pm$2.72 23.25$\pm$2.96
MOON-hetero Local 80.53$\pm$0.77 75.87$\pm$0.99 63.98$\pm$3.20 52.85$\pm$2.03 50.21$\pm$2.42 38.21$\pm$3.61 28.36$\pm$1.70 24.47$\pm$2.42 17.94$\pm$2.51
Global 88.05$\pm$0.14 87.92$\pm$0.17 79.05$\pm$0.89 58.39$\pm$0.20 57.55$\pm$0.17 55.01$\pm$0.34 31.28$\pm$1.61 30.68$\pm$1.92 24.91$\pm$1.78
Homo-FL FedFusion-hetero Local 82.35$\pm$0.51 75.27$\pm$1.00 62.20$\pm$5.37 51.23$\pm$1.73 48.21$\pm$2.08 37.53$\pm$5.15 33.98$\pm$2.07 25.62$\pm$2.29 20.31$\pm$5.25
Global 86.69$\pm$0.30 85.70$\pm$0.15 78.47$\pm$1.11 59.53$\pm$0.11 58.86$\pm$0.12 56.12$\pm$0.35 35.05$\pm$1.87 31.56$\pm$1.23 26.09$\pm$0.73
FedGen-hetero Local 82.29$\pm$0.62 76.01$\pm$0.97 62.74$\pm$3.17 51.99$\pm$3.67 47.80$\pm$1.12 39.79$\pm$6.02 34.04$\pm$2.93 25.37$\pm$2.56 21.77$\pm$6.89
Global 87.34$\pm$0.07 86.81$\pm$0.08 79.22$\pm$2.03 58.49$\pm$0.06 57.63$\pm$0.86 56.05$\pm$0.52 34.71$\pm$1.90 32.00$\pm$1.56 26.84$\pm$1.52
DaFKD-hetero Local 83.83$\pm$1.01 77.69$\pm$2.06 64.71$\pm$2.29 52.76$\pm$3.33 48.99$\pm$1.73 39.98$\pm$5.09 33.85$\pm$0.96 26.02$\pm$3.19 22.26$\pm$4.49
Global 89.26$\pm$0.31 87.84$\pm$0.64 80.07$\pm$1.23 58.97$\pm$0.43 58.52$\pm$0.19 57.33$\pm$1.00 35.69$\pm$1.48 32.57$\pm$1.34 26.54$\pm$1.73
Hetero-FL FedMD Local 70.33$\pm$3.98 63.47$\pm$4.13 60.67$\pm$5.09 42.28$\pm$6.11 39.64$\pm$5.45 26.90$\pm$7.70 23.84$\pm$3.76 19.37$\pm$3.20 13.89$\pm$6.70
Global 78.91$\pm$2.32 75.48$\pm$3.31 66.67$\pm$2.50 46.37$\pm$3.11 49.40$\pm$4.65 39.44$\pm$9.89 26.37$\pm$2.80 22.60$\pm$2.43 19.42$\pm$3.17
MSFKD Local 82.26$\pm$0.63 77.31$\pm$3.04 62.52$\pm$9.36 52.06$\pm$2.20 49.94$\pm$1.27 39.72$\pm$3.70 34.30$\pm$3.11 27.93$\pm$2.71 22.59$\pm$5.84
Global 87.79$\pm$0.18 86.98$\pm$0.31 80.00$\pm$1.75 58.88$\pm$0.38 59.09$\pm$0.24 56.29$\pm$0.83 36.29$\pm$0.60 31.62$\pm$0.87 26.70$\pm$2.33
FedGD Local 82.40$\pm$1.01 76.02$\pm$1.39 63.71$\pm$8.06 51.02$\pm$2.07 50.17$\pm$1.86 39.17$\pm$2.78 35.08$\pm$2.55 26.50$\pm$1.64 23.47$\pm$4.83
Global 87.52$\pm$0.24 87.22$\pm$0.13 79.31$\pm$0.75 59.26$\pm$0.41 58.03$\pm$0.26 56.34$\pm$0.65 36.86$\pm$1.06 30.66$\pm$1.59 27.53$\pm$2.20
FedFD (ours) Local 84.91$\pm$0.42 78.03$\pm$1.49 65.33$\pm$9.22 54.98$\pm$1.76 52.24$\pm$1.90 41.68$\pm$4.12 36.78$\pm$5.13 30.90$\pm$5.25 23.41$\pm$4.99
Global 90.06$\pm$0.03 89.64$\pm$0.23 82.74$\pm$0.58 61.07$\pm$0.22 60.86$\pm$0.10 59.24$\pm$0.48 40.27$\pm$1.33 34.24$\pm$1.13 29.09$\pm$1.69
@@ 4.2 Performance Overview
HeteroFL 12$\pm$3 25$\pm$6 20$\pm$3 $>200$ 124$\pm$16 $>200$
MOON-hetero 11$\pm$4 26$\pm$7 22$\pm$9 $>200$ 100$\pm$9 188$\pm$7
FedFusion-hetero 20$\pm$7 52$\pm$7 61$\pm$11 $>200$ 96$\pm$25 153$\pm$28
FedGen-hetero 21$\pm$11 38$\pm$15 198$\pm$20 $>200$ 115$\pm$12 176$\pm$24
DaFKD-hetero 23$\pm$9 41$\pm$9 44$\pm$13 191$\pm$27 85$\pm$8 126$\pm$21
FedMD 62$\pm$20 $>200$ 139$\pm$21 $>200$ $>200$ $>200$
MSFKD 19$\pm$9 45$\pm$17 33$\pm$10 171$\pm$20 87$\pm$22 115$\pm$31
FedGD 17$\pm$8 34$\pm$5 40$\pm$12 189$\pm$21 106$\pm$14 150$\pm$23
FedFD (ours) 10$\pm$5 20$\pm$9 16$\pm$6 60$\pm$18 67$\pm$8 99$\pm$19
@@
Ablation Study. As shown in Table 4.2, we evaluate the effects of each module in our model via ablation studies. -w/o feature alignment and -w/o orthogonal projection denote the performance of our model without using hierarchical feature alignment and orthogonal projection. Compared with FedFD, the performance of FedFD -w/o orthogonal projection and FedFD -w/o orthogonal projection degrades evidently with a range of 0.63$\sim$2.43%. While moving both components, the performance will drop significantly. Specifically, the orthogonal projection technique plays a much more important role with the obvious improvement in test accuracy, and the hierarchical feature alignment technique has a relatively small impact on test accuracy, which may be due to the scale of the experiment, but it can save storage costs. Experiment results verify that all modules are essential to train a robust Hetero-FL model.
@@ 5 Conclusion
Limitations. Although existing works and our method have demonstrated significant effectiveness in FD scenarios, none have investigated the case of LLMs. Nowadays, LLMs have emerged as the state-of-the-art technology in productivity, gradually replacing traditional small models. However, hardware limitations make deploying LLMs on client-side devices extremely challenging. Additionally, traditional knowledge distillation methods fail to address the knowledge transfer issue between small models on the client side and large models on the server side. In the future, we seek to move forward in this field.

### v2

@@ 1 Introduction
To address this issue, knowledge distillation DBLP:journals/corr/abs-1811-11479 has emerged as a promising approach, focusing on aggregating the output soft predictions of multiple local models into the global model. The authors in wu2022communication propose aggregating the logit output instead of model parameters and li2019fedmd utilize the aggregated class score to regulate the local training process. Building on the partial parameter aggregation proposed in diao2020heterofl , many works use the ensemble distillation to improve the global model like lin2020ensemble ; wang2023dafkd . wang2022knowledge and zhang2023target have focused on knowledge selection with logit distillation to optimize the aggregated model. shao2024selective identifies the accurate and precise knowledge from local and ensemble predictions.
@@ 3.3 FedFD: Feature Distillation for Hetero-FL
Instead of maintaining the projection layer separately for each client, the server now only needs to train (m1)(m-1) projection layers $\{\mathcal{M}_{2},\ldots,\mathcal{M}_{m}\}$. This not only reduces training parameters but also ensures that each projection layer has sufficient knowledge for distillation.
@@ 4.1 Experiment Setup
Configurations: Unless otherwise mentioned, we set the number of local training epoch E = 10, communication round T = 200, and the client number K = 20 with an active ratio r = 0.4. For local training, the batch size is 64 and the weight decay is 1e41e-4. The learning rate is $0.01$ for distillation and $0.001$ for training the local model. For the model on the server, we employ ResNet-18 he2016deep as the basic backbone. Like diao2020heterofl , we construct ten different computation complexity levels {a, b, c,…,j} with the hidden channel decay rate 10%. For example, model "a" has all the model parameters, while models "b" and "c" have the 90% and 80% effective parameters. We use "a-d-g" three different model architectures in our experiments and conduct more experiments with other architectures in Section 4.2. Similarly to diao2020heterofl , we have conducted statistics based on the performance of the global model for all secondary experiments.
@@
[…]
@@ 4.2 Performance Overview
[…]
@@
[…]